# An Investigation of the Current Star Formation Rate of Star-forming Galaxies at z < 0.5 with FADO

### v1

@@ Abstract
The star formation rate (SFR) is a crucial astrophysical tracer for understanding the formation and evolution of galaxies, determining the interaction between interstellar medium properties and star formation, thereby inferring the evolutionary laws of cosmic star formation history and cosmic energy density. The mainstream approach to study the stellar content in galaxies relies on pure stellar population synthesis models. However, these methods fail to account for the contamination of SFR caused by nebular gas radiation. Recent studies have indicated that neglecting nebular radiation contamination appears to be non-negligible in galaxies with intense star-forming activities and at relatively high redshifts, potentially leading to overestimation of stellar masses. However, there is currently limited targeted research, particularly regarding galaxies at higher redshifts (z 0.5). In this investigation, we employ the BPT diagram to select a sample of 2575 star-formation galaxies (SFG) from the SDSS-DR18 dataset, all within specified signal-to-noise ratio (S/N) ranges. Using the spectroscopic fitting tool FADO, which is capable of excluding nebular radiation contributions in spectral fitting, we conduct a tentative investigation of the SFR of star-forming galaxies in SDSS-DR18 with redshifts z 0.5. Our results show that 45 of the samples show H6563 obtained from FADO fitting to be smaller than that derived solely from the pure stellar population synthesis model qsofitmore, particularly pronounced between redshifts 0.2 and 0.4. After calculating the SFR from H6563 luminosity, we find that the contribution of nebulae is significant and exhibits an evolutionary trend with redshift, consistent with recent speculative findings. We anticipate that by combining optical and near-infrared spectral data, the influence of nebulae may become more prominent in star-forming galaxies at higher redshifts (e.g., up to z 2).
@@ 1 Introduction
The star formation rate (SFR), particularly the instantaneous SFR (SFR(t)), serves as a significant tracer for tracking stellar formation activities in galaxies, understanding their evolutionary history, and elucidating galaxy chemical evolution (Zhuang & Ho, 2019). It characterizes the rate at which new stars form within galaxies, thereby influencing their mass, type, and morphology. By combining observations of SFR with galaxy evolution simulations, the dominant physical processes driving galaxy growth can be thoroughly understood. For instance, processes such as galaxy quenching, gas accretion, star formation, and feedback from supernovae and active galactic nuclei (AGN) (Krumholz et al., 2012; Mullaney et al., 2012). Stellar formation represents the end point of a series of physical processes. Initially, gas cools from the interstellar medium and accretes onto the disk, forming a cold ‘atomic phase’, which then contracts into gravitationally bound clouds. Subsequently, molecular and molecular clouds form, dense clumps develop within molecular clouds, eventually leading to the formation of protostellar cores, stars, and star clusters. Although the starting and ending points of this process are relatively clear, the sequence of intermediate steps, especially the roles of forming cold atomic gas, molecular gas, and bound clouds, remains unclear, with different processes potentially dominating at different stages (Kennicutt & Evans, 2012). Investigating the influence of nebulae on SFR in galaxies contributes to inferring how gas affects the stellar formation process within galaxies and describing the conditions that compel gas to form stars. With increasingly reliable and precise estimations of SFR in galaxies and the discovery of numerous high-redshift galaxies by JWST, academic interest in researching SFR continues to soar. This has led to a burgeoning investigation into the relationship between redshift and SFR, sparking further inquiries in the field.
@@
SFR also holds significant implications for a broad range of astrophysical studies. The Logarithmic linear relationship between stellar mass and SFR, known as the star-forming main sequence (SFMS) (Noeske et al., 2007), aids in understanding the trajectory of stellar activities in the Hertzsprung-Russell (HR) diagram and analyzing the characteristics of stars during the main sequence phase (Tacconi et al., 2020). SFR is crucial for cosmological reionization studies. When the first batch of O-type stars (Pop III) formed in the universe, primarily composed of hydrogen and helium, their intense radiation led to cosmic reionization. Accurate SFR estimation is also vital for studying the cosmic microwave background (CMB) radiation. By measuring SFR at different epochs in cosmic history, the contribution of various types of galaxies to the cosmic background radiation can be estimated, thereby refining cosmological ionization models and searching for the Pop III in the universe (Madau & Dickinson, 2014). Therefore, relatively precise SFR measurements are beneficial for various properties in astrophysics.To date, numerous large-scale astrophysical projects have been related to SFR. For instance, variations in SFR during the formation of star clusters and massive stars in different environments (Lada & Lada, 2003; McKee & Tan, 2002; Kirk et al., 2015; McKee & Ostriker, 2007); the promotion of SFR or black hole growth due to interactions between galaxies (Di Matteo et al., 2005; Springel et al., 2005); SFR corrections for cosmological reionization (Madau & Dickinson, 2014; Wang, 2013; Robertson & Ellis, 2012; Daigne et al., 2004); reevaluation of star formation within individual molecular clouds and gas relations (Kennicutt & Evans, 2012). Several large-scale surveys in astrophysics also focus on studying the properties of star-forming galaxies, such as the Sloan Digital Sky Survey (SDSS) (Peng et al., 2010), the Hubble Ultra Deep Field (HUDF) (Ellis et al., 2013), and the Atacama Large Millimeter Array (ALMA) (Wang et al., 2013).
@@
Mainstream methods for estimating the SFR typically involve a fundamental step of spectral fitting for galaxies. This process entails fitting the spectrum as a combination of pure stellar population components, utilizing models based solely on stellar populations as templates for galaxy spectral fitting (e.g., STARLIGHT Cid Fernandes et al., 2005), Ppxf (Cappellari & Emsellem, 2004), qsofitmore (Fu, 2021). Considering that SDSS galaxies exhibit minimal nebular radiation due to insignificant star-forming activities, the aforementioned approaches overlook the modeling of nebular gas (Miranda et al., 2023), thereby neglecting the influence of nebular emission (Miranda et al., 2023; Ahumada et al., 2020). The limitation of these pure stellar models lies in their inability to solve different processes of gas emission within galaxies, as they incorporate assumptions that combine stellar mass, gas mass, and dust mass into a singular framework. However, studies by Krueger et al. (1995) and Salim et al. (2007) indicate that vigorous star-forming activities can lead to the nebular continuum contributing approximately 30%-70% to optical and near-infrared emissions, with nebular emission around star-forming regions accounting for about 60% of the total emission (Schaerer & de Barros, 2009; Krumholz et al., 2012). Therefore, when the acquired galaxy spectrum contain star-forming regions within the galaxy, the influence of nebulae becomes more pronounced. Pacifici et al. (2015) demonstrated that neglecting nebular emission in spectral modeling could lead to an overestimation of SFR by approximately 0.12 dex. Ignoring the contribution of nebular continuum would result in larger stellar mass and shallower ultraviolet slopes (Izotov et al., 2024). Galaxy evolution is primarily driven by the transformation of interstellar gas into stars, gas accretion between galaxies, intergalactic gas collisions, and galaxy mergers (Scoville et al., 2023). Thus, it is evident that interstellar gas plays a crucial role in the evolution of stars. FADO presents a viable solution for addressing the neglect of nebular emission in spectral processing. As the first independent and self-consistent spectral fitting tool to simultaneously consider stellar and nebular components in galaxies, FADO accounts for the impact of nebulae on SFR, dynamically and synchronously decomposing and calculating the contributions of different components and radiation mechanisms within the entire galaxy spectrum (Gomes & Papaderos, 2017). Therefore, in this investigation, we employ FADO as a research tool to consider nebular emission.
@@
In this survey, we selected spectra of galaxies with active star formation at relatively high redshifts (z 0.5) from SDSS-DR18. These spectra were fitted using the FADO software and the Python toolkit qsofitmore, based on stellar population synthesis models. The aim was to explore whether nebular emission would interfere with the intrinsic estimation of SFR in galaxies at higher redshifts. H6563 and H4861 are commonly regarded as the most reliable estimators of SFR, with Sobral et al. (2013) subsequently demonstrating the usefulness of H up to at least z 2.2. However, more extensive consideration of the [O II]3728 line is required around z 0.5, which can serve as an estimator of SFR up to z 1.4 (Kewley et al., 2004). SFR below z 0.4 can be accurately traced by H (Kennicutt, 1998), moderate redshift SFR by [O II]3728 (Kewley et al., 2004), and high redshift SFR by [O III]5007 (Shapley et al., 2023). Compared to the rough estimate of SFR provided by [O II]3728, H6563 entails fewer assumptions, smaller uncertainties, and more accurate constraints on star-forming regions, making it the most accurate tracer of SFR to date (Kennicutt & Evans, 2012). However, H6563 imposes strict constraints on redshift due to its longer wavelength, which may result in redshifts between 0.4 and 0.5 being beyond the observable wavelength range of the largest SDSS telescopes. Detailed analysis will be presented in Section 2.1. Furthermore, stellar age can be traced by the Balmer discontinuity at 3646 Å and the 4000 Å break; Si4128 and Cii1336 can track gas outflows or inflows; Mg2796 and [O iii]5007 can also estimate stellar metallicity and temperature (Förster Schreiber & Wuyts, 2020).
@@
Several studies have recently demonstrated the stability of FADO in spectral fitting and line information estimation. Breda et al. (2022) found, through comparison with STARLIGHT using simulated data, STARLIGHT overestimated the mass-weighted mean stellar age by up to 4 dex and the metallicity and light-weighted mean stellar age by up to 2 dex. Given these findings, they concluded that FADO can accurately recover mass, age, and mean metallicity with high precision ( 0.2 dex). Furthermore, Cardoso et al. (2022) showed that for galaxies with EW (H) >3 , the average mass and metallicity of the entire galaxy can differ by up to 0.06 dex, with a discrepancy of 0.12 dex. Most recently, Miranda et al. (2023) compared the SFR obtained from FADO fitting of the MPA-JHU galaxy sample with that obtained from STARLIGHT fitting of the MPA-JHU sample and found a difference of 0.01 dex in Ha flux between the two methods. They also predicted that nebular radiation not only plays a crucial role in the star formation activity of galaxies but also affects the calculation of galaxy SFR, particularly in high-redshift galaxies. In conclusion, the presence of nebulae does indeed affect the accurate estimation of SFR. Additionally, recent research by Izotov et al. (2024) has shown that neglecting the contribution of nebular continuum in determining SFR would result in larger stellar masses and shallower ultraviolet slopes. This is particularly significant for galaxies like J2229+2725, which are considered analogs of early-universe dwarf galaxies and play a role in cosmic reionization.
@@
The purpose of this investigation is to simultaneously fit galaxy spectra using the pure stellar code qsofitmore and FADO, which considers the contribution of nebular continuum. By examining the differences in the H6563 emission line in the fitted spectra, we aim to further determine the disparity in SFR. It is worth mentioning that based on the rough information available about stars in galaxies, stellar mass and SFR may not be sufficient to infer the potential evolutionary paths of galaxies. We still need to combine the spatial resolution information provided by all baryonic components of galaxies (stars, gas, metals) and the complete kinematic tracers of gravitational potential (dark matter), as well as the kinematic tracers of galaxy feedback processes (quenching, AGN, and supernova feedback), in order to obtain the complete evolutionary paths of galaxies (Förster Schreiber & Wuyts, 2020). Therefore, we remain open-minded regarding the derivation of galaxy evolution processes in the current scenario where we have no knowledge of dark matter.
@@
Section 2 outlines the criteria for selecting sources in the galaxy spectroscopic analysis and the process of spectral fitting employed in this study. In Section 3, we present the statistical analysis of the selected sample and the main findings. Section 4 discusses four critical issues related to the spectroscopic data and fitting techniques employed in this investigation. Finally, Section 5 provides a summary of this preliminary research. Throughout the paper, we assume a standard cosmological model with Hubble constant = 72 km s Mpc, matter density = 0.3, and cosmological constant = 0.7.
@@
The galaxy sample and their spectra used in this investigation are selected from Sloan Digital Sky Survey Data Release 18th (SDSS-DR18, Almeida et al., 2023).It integrates data from the previous 17 data releases and incorporates new observational data and improvements. Spectra were retrieved from the optical spectroscopic section of the SDSS Science Archive Server (SAS) 111https://dr18.sdss.org/optical/spectrum/search. We selected all galaxies with redshifts not exceeding 0.5, resulting in approximately 1.73 million samples.
2.1 Parent sample
To ensure reliable spectral fitting results, it is necessary to select samples with a sufficient S/N. Considering that FADO performs better in fitting spectra with median S/N > 5 compared to other pure stellar codes (Pappalardo et al., 2021), we first selected samples with a median S/N greater than 5. Furthermore, to effectively utilize high-quality spectra at arbitrary redshifts and ensure a uniform distribution of galaxy samples at different redshifts, we set an exponential curve (red solid line in Figure 1) and selected 36,109 galaxies with high S/N (blue scatter in Figure 1), constituting the parent sample. In fact, for the majority of spectroscopic surveys, there exists an inverse exponential relationship between the redshift of extragalactic sources and the maximum value of the spectroscopic S/N.As the redshift increases, the number of celestial objects captured by SDSS decreases(Almeida et al., 2023). Thus, this exponential curve is reasonable. The analytical expression of this curve is as follows:
From the histogram in the upper panel of Figure 1, it can be seen that the distribution of the parent sample’s quantity versus redshift is relatively uniform. This ensures that our conclusions are not excessively affected by sample bias. However, as will be seen in the following two sections, further processing of the parent sample and stringent selection criteria inevitably result in the loss of a large number of samples, and even cause the extracted new samples to no longer be uniformly distributed with redshift. Nevertheless,the parent sample have to be addressed in depth to obtain a pure sample, that is, the gold sample. The parent sample have to be addressed in depth to obtain a pure sample.
@@ 2.2 BPT selection
The data selection in this survey involves classifying galaxies based on their positions on the BPT diagram. We utilize the BPT diagram based on four emission line ratios, namely [N II] 6583, [O III] 5007, 6563, and 4861 (Baldwin et al., 1981). The selected data points are ensured to lie in the pure star-forming (SF) region of the BPT diagram, with S/N 5 and z 0.5. The two boundary lines in Figure 2 are derived from Equation 2 (Kewley et al., 2001) and Equation 3 (Kauffmann et al., 2003). These boundary lines divide the BPT diagram into three regions, where the blue represents star-formation galaxies, gray represents composite galaxies, and red represents AGN regions. Prior to inclusion in the BPT diagram, we carefully select galaxies from 36,109 with reasonable flux values for the aforementioned four emission lines, ensuring that these data points have physically meaningful positive fluxes. Subsequently, the galaxies selected as SF galaxies (SFG) are plotted on the BPT diagram to obtain the required SFG distribution, as depicted in Figure 2. A total of 2575 galaxies are ultimately identified as SFGs on the BPT diagram. These data points are then subjected to spectral fitting using qsofitmore and FADO to obtain the most accurate emission line fluxes.

### v3

# Exploring the Current Star Formation Rate and Nebula Ratio of Star-Formation Galaxies at z < 0.4 with FADO

@@ Abstract
The star formation rate is a crucial astrophysical tracer for understanding the formation and evolution of galaxies, determining the interaction between interstellar medium properties and star formation, thereby inferring the evolutionary laws of cosmic star formation history and cosmic energy density. The mainstream approach to studying the stellar property in galaxies relies on pure stellar population synthesis models. However, these methods fail to account for the contamination of SFR caused by nebular gas radiation. Recent studies have indicated that neglecting nebular radiation contamination appears non-negligible in galaxies with intense star-forming activities and at relatively high redshifts, potentially leading to overestimating stellar masses. However, there is currently limited targeted research, particularly regarding galaxies at redshifts (z 0.4). In this work, 6,511 star-formation galaxies are selected from the SDSS-DR18, and FADO fits their spectra. This tool can exclude nebular radiation contributions in the spectral fitting. A tentative work is carried out to explore the SFR of these galaxies. The results indicate that the median flux obtained from FADO fitting differs from that obtained using the pure stellar population synthesis model qsofitmore by approximately 0.034 dex. Preliminary evidence suggests that the average nebula ratio increases with redshift. Additionally, we investigated the impact of stellar mass on the nebula ratio at low to moderate redshifts. By comparing two spectral fitting software packages, we found that although the contribution of nebular emission is minimal, it generally shows an increasing trend with redshift. We anticipate that by combining optical and near-infrared spectral data, the influence of nebulae may become more prominent in star-forming galaxies at higher redshifts (e.g., up to z 2).
@@ 1 Introduction
The star formation rate (SFR), particularly the instantaneous SFR (SFR(t)), serves as a significant tracer for tracking stellar formation activities in galaxies, understanding their evolutionary history, and elucidating galaxy chemical evolution (Zhuang & Ho, 2019). By combining observations of SFR with galaxy evolution simulations, the dominant physical processes driving galaxy growth can be thoroughly understood. For instance, processes such as galaxy quenching, gas accretion, star formation, and feedback from supernovae and active galactic nuclei (AGN) (Krumholz et al., 2012; Mullaney et al., 2012). The logarithmic linear relationship between stellar mass and SFR, known as the star-forming main sequence (SFMS, e.g., Noeske et al., 2007), aids in understanding the trajectory of stellar activities in the Hertzsprung-Russell (HR) diagram and analyzing the characteristics of stars during the main sequence phase (Tacconi et al., 2020). By measuring SFR at different epochs in cosmic history, the contribution of various types of galaxies to the cosmic background radiation can be estimated, thereby refining cosmological ionization models and searching for the Pop III in the universe (Madau & Dickinson, 2014). To date, numerous efforts have been put into the SFR, for instance, variations in SFR during the formation of star clusters and massive stars in different environments(Lada & Lada, 2003; McKee & Tan, 2002; Kirk et al., 2015; McKee & Ostriker, 2007); the promotion of SFR or black hole growth due to interactions between galaxies (Di Matteo et al., 2005; Springel et al., 2005); SFR corrections for cosmological reionization (Madau & Dickinson, 2014; Wang, 2013; Robertson & Ellis, 2012; Daigne et al., 2004); reevaluation of star formation within individual molecular clouds and gas relations (Kennicutt & Evans, 2012). Several surveys in astrophysics involved studying the properties of star-forming galaxies, such as the Sloan Digital Sky Survey (SDSS, York et al., 2000), the Hubble Ultra Deep Field (HUDF, Ellis et al., 2013), and the Atacama Large Millimeter Array (ALMA, Wang et al., 2013).
@@
Mainstream methods for estimating the SFR typically involve a fundamental step of spectral fitting for galaxies. This process entails fitting the spectrum as a combination of pure stellar population components, utilizing models based solely on stellar populations as templates for galaxy spectral fitting (e.g., STARLIGHT (Cid Fernandes et al., 2005), pPXF (Cappellari & Emsellem, 2004), qsofitmore (Fu, 2021)) Considering that SDSS galaxies exhibit minimal nebular radiation due to insignificant star-forming activities, the approaches above overlook the modelling of nebular gas (Miranda et al., 2023), thereby neglecting the influence of nebular emission (Miranda et al., 2023; Ahumada et al., 2020). The limitation of these pure stellar models lies in their inability to solve different processes of gas emission within galaxies, as they incorporate assumptions that combine stellar mass, gas mass, and dust mass into a singular framework. However, studies by Krueger et al. (1995) and Salim et al. (2007) indicate that vigorous star-forming activities can lead to the nebular continuum contributing approximately 30%-70% to optical and near-infrared emissions, with nebular emission around star-forming regions accounting for about 60% of the total emissions (Schaerer & de Barros, 2009; Krumholz et al., 2012). Therefore, when the acquired galaxy spectrum contains star-forming regions within the galaxy, the influence of nebulae becomes more pronounced. Pacifici et al. (2015) demonstrated that neglecting nebular emission in spectral modelling could lead to an overestimated SFR by approximately 0.12 dex. Ignoring the contribution of the nebular continuum would result in larger stellar mass and shallower ultraviolet slopes (Izotov et al., 2024). Galaxy evolution is primarily driven by transforming interstellar gas into stars, gas accretion between galaxies, intergalactic gas collisions, and galaxy mergers (Scoville et al., 2023). As the first independent and self-consistent spectral fitting tool to simultaneously consider stellar and nebular components in galaxies, FADO accounts for the impact of nebulae on SFR, dynamically and synchronously decomposing and calculating the contributions of different components and radiation mechanisms within the entire galaxy spectrum (Gomes & Papaderos, 2017). Therefore, we employ FADO as a research tool to consider nebular emission in this investigation.
@@
In this work, we selected spectra of galaxies with active star formation at middle redshifts (z 0.4) from SDSS-DR18. These spectra were fitted using the FADO software and the Python toolkit qsofitmore, based on stellar population synthesis models. The aim was to explore whether nebular emission would interfere with the intrinsic estimation of SFR in galaxies at higher redshifts. and are commonly regarded as the most reliable estimators of SFR (Sobral et al., 2013). However, more extensive consideration of the [O II]3728 line is required around z 0.5, which can serve as an estimator of SFR up to z 1.4 (Kewley et al., 2004). SFR below z 0.5 can be accurately traced by (Kennicutt, 1998), moderate redshift SFR by [O II]3728 (Kewley et al., 2004), and high redshift SFR by [O III]5007 (Shapley et al., 2023).
@@
Several studies have recently demonstrated the stability of FADO in spectral fitting and line information estimation. Breda et al. (2022) found, through comparison with STARLIGHT using simulated data, STARLIGHT overestimated the mass-weighted mean stellar age, light-weighted mean stellar age, and metallicity. These findings concluded that FADO can accurately recover mass, age, and mean metallicity with high precision ( 0.2 dex). Furthermore, Cardoso et al. (2022) showed that for galaxies with EW (Hα) >3 , the average mass and metallicity of the entire galaxy can differ by up to 0.06 dex. Most recently, Miranda et al. (2023) compared the SFR obtained from FADO fitting of the MPA-JHU galaxy sample with that obtained from STARLIGHT fitting of the MPA-JHU sample and found a difference of 0.01 dex in Ha flux between the two methods. They also predicted that nebular radiation not only plays a crucial role in the star formation activity of galaxies but also affects the calculation of galaxy SFR, particularly in high-redshift galaxies. Additionally, recent research by Izotov et al. (2024) has shown that neglecting the contribution of the nebular continuum in determining SFR would result in larger stellar masses and shallower ultraviolet slopes. This is particularly significant for galaxies like J2229+2725, which are considered analogues of early-universe dwarf galaxies and play a role in cosmic reionization.
@@
This work aims to compare the spectroscopic fitting of the pure stellar code qsofitmore and FADO to explore the possible impact of the nebular contribution on SFR estimation. And explore the relationship between this effect and redshift. By examining the differences in the emission line in the fitted spectra, we aim to further determine the disparity in SFR. It is worth mentioning that based on the rough information available about stars in galaxies, stellar mass and SFR may not be sufficient to infer the potential evolutionary paths of galaxies. We still need to combine the spatial resolution information provided by all baryonic components of galaxies (stars, gas, metals) and the complete kinematic tracers of gravitational potential (dark matter), as well as the kinematic tracers of galaxy feedback processes (quenching, AGN, and supernova feedback), to obtain the complete evolutionary paths of galaxies (Förster Schreiber & Wuyts, 2020). Therefore, we remain open-minded regarding the derivation of galaxy evolution processes in the current scenario.
@@
Section 2 outlines the criteria for selecting sources in the galaxy spectroscopic analysis and the process of spectral fitting employed in this study. In Section 3, we present the statistical analysis of the selected sample and the main findings. Section 4 discusses four critical issues related to the spectroscopic data and fitting techniques employed in this investigation. Finally, Section 5 summarizes this preliminary research. Throughout the paper, we assume a standard cosmological model with Hubble constant = 72 km s-1 Mpc-1, matter density = 0.3, and cosmological constant = 0.7.
@@
The galaxy sample and their spectra used in this work are selected from Sloan Digital Sky Survey Data Release 18th (SDSS-DR18; Almeida et al., 2023). It integrates data from the previous 17 data releases and incorporates new observations and improvements. The galaxy sample was obtained by SQL query, and their spectroscopic data were retrieved from the SDSS Science Archive Server (SAS) 111https://dr18.sdss.org/optical/spectrum/search.
2.1 SQL query
To guarantee the subsequent spectral fitting, it is necessary to select spectra with sufficient signal-to-noise ratio (S/N). Since FADO performs better with median S/N of spectra larger than 5 (Pappalardo et al., 2021), the SQL query was adopted to select high S/N spectra. The query obtained 91,252 spectra with 5 < S/N < 20 for 0 < z < 0.2 galaxies, and 6,604 with S/N > 5 for 0.2 < z < 0.4 galaxies. Additionally, the S/N of major emission lines of each spectrum was also constrained by SQL query with S/N > 5. The following conditions were added in the query to ensure the reliability of redshift measurements: zwarning = 0, veldisperr > 0, veldisp > 3 * veldisperr, and 80 < veldisp < 350 (see Appendix A for detailed SQL query).
@@ 2.2 BPT selection
After SQL selection, the BPT diagram (Baldwin et al., 1981) was adopted to isolate the SFGs from the sample. The two line ratios, [N II]6583/ and [O III]5007/, were selected to build the BPT diagram. Two boundaries divide the sample into three regions: star-forming, composite, and AGN. In Figure 1, the dashed boundary Equation (1) is derived from Kewley et al. (2001), while the dotted one Equation (2) from Kauffmann et al. (2003). The samples in the pure star-forming region of the BPT diagram were selected as the SFGs (blue scatter in Figure 1).
A total of 21,455 galaxies were considered SFGs, which is a composite of 18,176 galaxies at 0 < z < 0.2 (lower-redshift sample) and 3,279 galaxies at 0.2 < z < 0.4 (higher-redshift sample). To avoid the bias of the larger number of lower-redshift galaxies, 3,279 of them were randomly selected to stand for the lower-redshift sample. Ultimately, a sample containing 6,558 galaxies was submitted to fit by qsofitmore and FADO, aiming at deriving more accurate emission line fluxes. Figure 2 shows the distribution of the two samples with different redshift intervals.